OpenStudy (anonymous):

How do you evaluate log35 (8) using the change of base formula? Thanks!

OpenStudy (whpalmer4):

Use the change of base formula: \[\log_a x = \frac{\log_b x}{\log_b a}\] So if you have a calculator that will do common logarithms, usually written \(\log\) but meaning \(\log_{10}\), you could do \[\log_{35} x = \frac{\log_{10} x}{\log_{10} 35}\]